Part A A box of oranges which weighs 93 N is being pushed across a horizontal floor. As it moves, it is slowing at a constant rate of 0.90 m/s each second. The push force has a horizontal component of 20 N and a vertical component of 25 N downward. Calculate the magnitude of the frictional force between the box and floor. Express your answer using two significant figures. Hints Submit My Answers Give Up Part B Calculate the magnitude of the upward normal force of the floor on the box of oranges Express your answer using two significant figures. 图?
Answer #1

mass of the box = 93/9.8 = 9.5 kg


acceleration ax = 0.9 m/s^2

along horizontal Fnet = Fx - fx = m*ax

fx = Fx - m*ax


frictional force fx = 20-(9.5*0.9) = 11.45 N <---ANSWER

=====================


part (b)

along vertical


n - Fy - W = 0

normal force n = W + Fy = 93 + 25 = 118 N <<---ANSWER